# An Overview of 7726 User Reports: Uncovering SMS Scams and Scammer Strategies

### v1

@@ VI Related Work
Our work fits into the broader literature on suspicious SMS messages. We divide them by data source. Previous researchers have investigated public SMS gateways to understand the SMS ecosystem and identify SMS phishing text messages. While Reaves et al. [90, 91] and Moreno et al. [92] mention a minority of SMS messages containing malicious URLs, but they report a wide use of URL shorteners that cannot resolve. Instead, by enriching our dataset with telemetry captured as SMSs are reported, we gain the ability to, for instance, resolve these shorteners and reliably query relevant contextual information. Nahapetyan et al. identify over 67.9kk SMS phishing messages using identifiers such as phone numbers, email addresses, and One-Time Codes without removing spam [32].
Some have collated victim reports [93, 31] or crowdsourced them [94, 95]. Others used intensive techniques like in-depth interviews [20, 96] and analyzing news articles [97]. More relatedly, some researchers investigated SMS messages caught by XDRs and interacted with the scammers behind the messages, albeit for only one type of scam (Hi Mum/Dad) [30]. Similarly, Agarwal et al. investigated two months of blocked SMS text messages from one UK mobile network operator and manually categorized them into six scam types [19]. Our methods build on their insights into attacker strategies, but we see that just focusing on these six scam types offers an incomplete picture of the landscape.
It is common to mix together spam and scam messages. A few groups collect spam and smishing (scam) text messages reported by users on Twitter, but treat them all as spam [31, 98]. Another recent work crowdsources suspicious text messages from users in the US, categorizing them all as scams [33]. Srinivasan et al. consider the SMS reports users submit to the Federal Trade Commission (FTC) and third parties as spam [36]. Contrastingly, we break down this difference which we show to materially change our results.
Previous work has devised machine learning models to detect smishing messages [99, 100, 101, 102, 103, 104, 105, 106, 107], using old spam datasets [108, 95, 109]. We provide a methodological framework to differentiate spam and identify new scam types from user reports that could help researchers enhance these models.
@@ VII Conclusion
It is mandatory in the UK to reimburse fraud victims after they fall victim to a scam and directly send the fraudster money [110]. Because of this, UK infrastructure operators are increasingly being pressured by banks to reduce fraud. These incentives have played out in our work investigating hundreds of SMS scams reported during a 4 month period in 2024. We winnow down over 1mm user reports into 213kk scam texts using a careful, multi-layered methodology. We investigate these scam texts and find that most that get through the mobile network operators’s XDR have no URL. We hypothesize this is from direct pressure and vast industry experience blocking URL-based scams. We highlight the emerging trend of call back scams, where scammers entice victims to call them on a phone number. We encourage those designing filtering products like XDRs for mobile network operators to more carefully consider conversation scams, since they often lead to fraud, even if not from the start.

### v2

@@ VI Related Work
Our work fits into the broader literature on suspicious SMS messages. We divide them by data source. Previous researchers have investigated public SMS gateways to understand the SMS ecosystem and identify SMS phishing text messages. While Reaves et al. [reaves2016sending, reaves2018characterizing] and Moreno et al. [moreno2023your] mention a minority of SMS messages containing malicious URLs, but they report a wide use of URL shorteners that cannot resolve. Instead, by enriching our dataset with telemetry captured as SMSs are reported, we gain the ability to, for instance, resolve these shorteners and reliably query relevant contextual information. Nahapetyan et al. identify over 67.9kk SMS phishing messages using identifiers such as phone numbers, email addresses, and One-Time Codes without removing spam [nahapetyan2024sms].
Some have collated victim reports [twitterreports_eCrime, tang:ccs22] or crowdsourced them [timko2023commercial, chen_creating_2012]. Others used intensive techniques like in-depth interviews [lutforsmishing_interview, lorrie_phishingsoups] and analyzing news articles [Oestnews_ecrime]. More relatedly, some researchers investigated SMS messages caught by XDRs and interacted with the scammers behind the messages, albeit for only one type of scam (Hi Mum/Dad) [agarwal_usenix25]. Similarly, Agarwal et al. investigated two months of blocked SMS text messages from one UK mobile network operator and manually categorized them into six scam types [agarwal_imc24]. Our methods build on their insights into attacker strategies, but we see that just focusing on these six scam types offers an incomplete picture of the landscape.
It is common to mix together spam and scam messages. A few groups collect spam and smishing (scam) text messages reported by users on Twitter, but treat them all as spam [tang:ccs22, grier:ccs10]. Another recent work crowdsources suspicious text messages from users in the US, categorizing them all as scams [smishtank_timko24]. Srinivasan et al. consider the SMS reports users submit to the Federal Trade Commission (FTC) and third parties as spam [srinivasan2016understanding]. Contrastingly, we break down this difference which we show to materially change our results.
Previous work has devised machine learning models to detect smishing messages [hosseinpour2025poster, jain2020novel, jain2019feature, goel2018smishing, jain2018rule, joo2017s, mishra2020smishing, mishra2019content, sonowal2018smidca], using old spam datasets [tiago_spam, chen_creating_2012, delany2012sms]. We provide a methodological framework to differentiate spam and identify new scam types from user reports that could help researchers enhance these models.
@@ VII Conclusion
It is mandatory in the UK to reimburse fraud victims after they fall victim to a scam and directly send the fraudster money [appfraudreimburse]. Because of this, UK infrastructure operators are increasingly being pressured by banks to reduce fraud. These incentives have played out in our work investigating hundreds of SMS scams reported during a 4 month period in 2024. We winnow down over 1mm user reports into 213kk scam texts using a careful, multi-layered methodology. We investigate these scam texts and find that most that get through the mobile network operators’s XDR have no URL. We hypothesize this is from direct pressure and vast industry experience blocking URL-based scams. We highlight the emerging trend of call back scams, where scammers entice victims to call them on a phone number. We encourage those designing filtering products like XDRs for mobile network operators to more carefully consider conversation scams, since they often lead to fraud, even if not from the start.
@@ Acknowledgments
-A Open AI Prompt: Message Classification
You will receive a json object with an ‘id’ and a ‘message’. The ‘id’ is the id of the message and the ‘message’ is text of a SMS. Based on the instructions below, process the message and return a json object. Instructions: 1. Identify the brand or organization that the message is trying to impersonate in the text. Return empty if none. (”named_entity” key in the json object. This key should always be returned in the json.) 2. Classify the type of smishing message (”scam_type” key in the json. This key should always be returned in the json.) The scam_types can be: ”Hey mum/dad” - text addressed to mum/mom or a dad and asking to text/call back potentially giving a reason about using a different mobile number. ”Delivery/Parcel” - text impersonating a parcel/delivery company asking to click on a link, text back or call on a number ”Banking” - text impersonating a bank or a financial institution asking to click on a link, text back or call on a number ”Government” - text impersonating a government organization asking to click on a link, text back or call on a number ”Telecom” - text impersonating a mobile network operator asking to click on a link, text back or call on a number ”Wrong number” - text with a normal greeting or asking about someone or to reply back ”Spam” - illicit marketing message including casino, betting, random draws, etc ”Others” - If it is does not fit as one of the above category 3. If the ”scam_type” classified is ”Others”, then identify a category for the text. Return empty otherwise. (”other_category” key in the json object. This key should always be returned in the json.) 4. Every json object should include the ”id” of the message being classified.
-B Open AI Prompt: Detecting Scam Lures
You will receive a json object with an ‘id’ and a ‘message’. The ‘id’ is the id of the message and the ‘message’ is text of a scam SMS. Based on the instructions below, process the message and return a json object. Instructions: 1. Provide which lure principles apply for each text message (”lure_principles” key should be a list and always be provided in the json object. If you cannot detect any lure principles, leave the list empty.) Lure principles are: a) Distraction Principle - providing various reasons to distract the user. b) Authority Principle - providing trust to the user to not question authority. could be done by making references to legitimate entities. c) Herd Principle - encouraging a user to not miss out on opportunities by relating to the popularity of a scheme. convincing how others have won things or take the same risk. d) Dishonesty Principle - inviting users willingly and knowingly to participate in a fraudulent scheme. e) Kindness Principle - Fraudsters leverage the willingness of people to help others. f) Need and Greed Principle - leveraging user’s greed and offering attractive (monetary) benefits so user would take an action asked in the text. g) Time/Urgency Principle - putting time pressure on user so they make a rushed decision. 2. Every json object should include the ”id” of the message being classified.
-C Other Scam Types Examples
An Enforcement Agent has been scheduled to attend your property, call <<name>> NOW ON <<phone number>> to prevent this action. Quote ref <<reference number>> Can you do this EPC? <<postcode>> Requested day/time: Call to confirm Your fee: £40 Accept job here: [URL] <<First Name>> is due a Kennel Cough Vaccine on 20/4/24. Call us on <<phone number>> to book. Thank you. Ensure we hold the correct student finance information for your April payment by visiting:[URL] <<customer name>>, Customer ID: <<id>>. You should receive a letter from us soon regarding <<brand name>>. It is important that you resolve this matter online as soon as possible by visiting [URL] Thanks, <<brand name>>. [URL] You have not yet paid £39.39 due to <<brand name>> INSURANCE SERVICES GR. Got a question about your account? Text us on <<phone number>>. Ref JCF000052164
-D The Gunning Fog Index
Fog Index Reading level by grade
17 College graduate
16 College senior
15 College junior
14 College sophomore
13 College freshman
12 High school senior
11 High school junior
10 High school sophomore
9 High school freshman
8 Eighth grade
7 Seventh grade
6 Sixth grade
5 Fifth grade
TABLE XIII: Education level required as per the Gunning fog index.
Figure 8: Gunning Fog Index for six scam types with and without URLs.
Scam Type Banking Government Delivery/Parcel Telecom Wrong Number Hi Mum/Dad
Banking 6,327
Government 32 1,222
Delivery/Parcel 183 52 11,648
Telecom 47 88 57 5,742
Wrong Number 47 13 29 35 18,404
Hi Mum/Dad 51 0 52 10 490 8,748
TABLE XIV: Common Sender IDs scammers abuse to send multiple scams.